# A complete proof of the geometric Bombieri–Lang conjecture for ramified covers of abelian varieties

### v1

@@ 3.3 A positivity lemma
Since $)(f\circ\pi)({\rm Supp}(\widetilde{{\mathcal{E}}}$)) does not contain ${\mathcal{A}}\times\{t_{\infty}\}$, we can choose $\mathcal{E}$$^{\prime}$ such that ${\rm Supp}({\mathcal{E}}^{\prime})$ does not contain ${\mathcal{A}}\times\{t_{\infty}\}$.
@@
Recall that $\widetilde{{\mathcal{R}}}_{0}$ is the union of all irreducible components of $\widetilde{{\mathcal{R}}}$ which dominate some irreducible component of $\mathcal{R}$}. View $\widetilde{{\mathcal{R}}}_{0}$ as a $\mathbb{Q}$-Cartier divisor on $\widetilde{{\mathcal{X}}}$ such that for any irreducible component of $\widetilde{{\mathcal{R}}}_{0}$, its multiplicity in $\widetilde{{\mathcal{R}}}_{0}$ is exactly the same as its multiplicity in $\widetilde{{\mathcal{R}}}$. Denote $\widetilde{{\mathcal{R}}}_{1}=\widetilde{{\mathcal{R}}}-\widetilde{{\mathcal{R}}}_{0}$. Then $\widetilde{{\mathcal{R}}}_{1}$ is an effective $\mathbb{Q}$-Cartier divisor on $\widetilde{{\mathcal{X}}}$ whose image in $\mathcal{X}$} has codimension greater than 1.
@@
Let ${\mathcal{Z}}_{1}$ be the image of $\widetilde{{\mathcal{R}}}_{1}$ in ${\mathcal{A}}\times_{\mathbb{C}}T$. Use Raynaud–Gruson’s flattening theorem (cf. [RG, Théorème 5.2.2]) to the composition ${\mathcal{Z}}_{1}\hookrightarrow{\mathcal{A}}\times_{\mathbb{C}}T\to T$, we get a proper birational map $g\colon T^{\prime}\to T$ such that the strict transform ${\mathcal{Z}}^{\prime}_{1}$ of ${\mathcal{Z}}_{1}$ under $g$ is flat over $T^{\prime}$. Note that ${\mathcal{Z}}^{\prime}_{1}$ is a closed subscheme of ${\mathcal{A}}\times_{\mathbb{C}}T^{\prime}$, and the inverse image of ${\mathcal{Z}}_{1}$ in ${\mathcal{A}}\times_{\mathbb{C}}T^{\prime}$ is the union of ${\mathcal{Z}}^{\prime}_{1}$ and some $T^{\prime}$-vertical closed subscheme ${\mathcal{V}}_{T^{\prime}}$ of ${\mathcal{A}}\times_{\mathbb{C}}T^{\prime}$ (i.e. the image of ${\mathcal{V}}_{T^{\prime}}$ in $T^{\prime}$ is not Zariski dense). After removing finitely many points, we can assume that $t_{n}$ has a unique pre-image $t^{\prime}_{n}$ in $T^{\prime}$ which does not lie in ${\rm pr}_{T^{\prime}}({\mathcal{V}}_{T^{\prime}})$ for all $n\geq 1$. Furthermore, we can assume $t^{\prime}_{n}$ tends to $t^{\prime}_{\infty}$ in the Euclidean topology. By our definition of $\widetilde{{\mathcal{R}}}_{1},{\mathcal{Z}}_{1},{\mathcal{Z}}^{\prime}_{1}$, we see that for any $t^{\prime}\in T^{\prime}$, $({\mathcal{Z}}^{\prime}_{1})_{t^{\prime}}$ has codimension greater than 1 in ${\mathcal{A}}\times\{t^{\prime}\}$. Let $Z_{1}$ be the image of $({\mathcal{Z}}^{\prime}_{1})_{t^{\prime}_{\infty}}\subseteq{\mathcal{A}}\times\{t^{\prime}_{\infty}\}$ in ${\mathcal{A}}\times\{t_{\infty}\}$.
@@
Use Lemma 3.7 for the morphism ${\mathcal{D}}_{{\mathrm{red}}}\hookrightarrow{\mathcal{A}}\times_{\mathbb{C}}T\to T$ and the sequence $t_{n}\to t_{\infty}$ of points of $T$, we get a Zariski open dense subset ${\mathcal{D}}_{t_{\infty}}^{\circ}\subseteq{\mathcal{D}}_{t_{\infty}}$ satisfying the properties in Lemma 3.7. Furthermore, taking a subsequence of $\{t_{n}\}_{n\geq 1}$ if necessary, we may assume that $\{t_{n}\}_{n\geq 1}$ satisfies the conditions of remark 3.9. Denote ${\mathcal{D}}_{b,t_{\infty}}^{\circ}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}({\mathcal{D}}_{t_{\infty}}^{\circ})_{b}$. We will essentially use the property of ${\mathcal{D}}_{t_{\infty}}^{\circ}$ in Sub-lemma 3.20 below, but for the sake of logical sequence, we need to choose this ${\mathcal{D}}_{t_{\infty}}^{\circ}$ at first.
@@
  2. 2.
    $S_{0}$ contains all the points $b\in B(\mathbb{C})$ such that $({\mathcal{D}}_{\rm red})_{b}$ or $(\widetilde{{\mathcal{R}}}_{\rm red})_{b}$ is non-reduced, which is finite by the generic reducedness (cf. [Sta, Lemma 054Z, Lemma 020I, Lemma 0578]).
  3. 3.
@@
Take $\ell_{n}\to\infty$ such that $\ell^{-1}_{n}s_{n}$ converges to a nonzero element $s_{\infty}$ in $V(A,K)$. Let $G=G(s_{\infty})$ be the smallest abelian subvariety of $A$ such that $s_{\infty}$ lies in $G(K)_{\mathbb{R}}$. Denote by $\mathcal{G}$} the Zariski closure of $G$ in $\mathcal{A}$}.
@@
As $b\notin S_{0}$, by property (6) of $S_{0}$, we can take an open disk ${\mathbb{D}}\subseteq B(\mathbb{C})$ with center $b$ such that $\mathbb{D}$} does not intersect with the image of ${\rm Supp}(\widetilde{{\mathcal{V}}}_{B})$ in $B$. After removing finitely many $n$, we may assume that $b_{n}\in{\mathbb{D}}$ for all $n\geq 1$. Apply Theorem 2.4 to $\{s_{n}\}_{n\geq 1}$ and the limit point $x\in({\mathcal{A}}_{b}\times\{t_{\infty}\})\backslash({\mathcal{D}}_{b,t_{\infty}}\cup{\mathcal{E}}^{\prime}_{b,t_{\infty}}\cup(Z_{1})_{b})$ of $\{s_{n}(B)\}_{n\geq 1}$. We obtain a linear entire curve $\varphi_{\infty}=\phi_{(x,\delta(s_{\infty},v_{b}))}\times\{t_{\infty}\}\colon\mathbb{C}\to{\mathcal{A}}_{b}\times\{t_{\infty}\}$, which is the unique limit of the re-parametrization $\{\varphi_{n}=\phi_{n}\times\{t_{n}\}\colon{\mathbb{D}}_{r_{n}}\to{\mathcal{A}}\times\{t_{n}\}\}_{n\geq 1}$ of $\{(s_{n},t_{n})\colon B\to{\mathcal{A}}\times\{t_{n}\}\}_{n\geq 1}$. Explicitly, recall that $\varphi_{n}$ is defined by
@@
For a line bundle $L$ on ${\mathcal{A}}\times_{\mathbb{C}}T$ belongs to one of the above (which has 2 possibilities) with the fixed metric $\left\|\cdot\right\|_{L}$, we choose the induced metric $(f\circ\pi)^{*}\left\|\cdot\right\|_{L}$ on $(f\circ\pi)^{*}L$. Moreover, for any $\mathbb{Z}$-linear combination of the line bundles above (including the two pullback line bundles), we can also fix a natural metric on this new line bundle induced by the metrics above.
For a line bundle $L$ which is a $\mathbb{Z}$-linear combination of the line bundles above (including the three pullback line bundles) with the fixed metric $\left\|\cdot\right\|_{L}$, a holomorphic map $\phi\colon{\mathbb{D}}_{R}\to\widetilde{{\mathcal{X}}}$ (or $\phi\colon{\mathbb{D}}_{R}\to{\mathcal{A}}\times_{\mathbb{C}}T$, depending on $L$ is a line bundle on $\widetilde{{\mathcal{X}}}$ or on ${\mathcal{A}}\times_{\mathbb{C}}T$), set
@@
Finally, the above definitions can be extended by $\mathbb{Q}$-linearity to any $\mathbb{Q}$-linear combination of the line bundles above (including the three pullback line bundles).
@@
By assumption (2) for $x$, $({\mathcal{D}}_{b,t_{\infty}})_{\rm red}\cap(x+{\mathcal{G}}_{b}\times\{t_{\infty}\})$ is a reduced divisor on $(x+{\mathcal{G}}_{b}\times\{t_{\infty}\})$. Using Theorem 3.6(1) and the ampleness of ${\mathcal{D}}_{b,t_{\infty}}$ on ${\mathcal{A}}_{b}\times\{t_{\infty}\}$ , we see that there exists a subset $\Sigma_{\varepsilon}\subseteq\mathbb{R}^{+}$ of finite Lebesgue measure such that for any $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$,
@@
Therefore for any fixed $\varepsilon>0$ and any $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$, combine (3.4.4) and (3.4.6), we get
@@
Moreover, for fix $0<\varepsilon<\frac{1}{2}$ above, $\frac{\varepsilon}{1-\varepsilon}<2\varepsilon$ and $N^{(1)}(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})\geq\frac{c_{2}}{\varepsilon}$ when $r$ large enough. Thus by enlarging $\Sigma_{\varepsilon}$ we may assume that for all $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$,
@@
On the other hand, combine (3.4.4) and (3.4.5), we have for all $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$,

### v2

@@ 3.4 Step 2: lower bound of transversal intersections
Since $)(f\circ\pi)({\rm Supp}(\widetilde{{\mathcal{E}}}$)) does not contain ${\mathcal{A}}\times\{t_{\infty}\}$, we can choose $\mathcal{E}$$^{\prime}$ such that ${\rm Supp}({\mathcal{E}}^{\prime})$ does not contain ${\mathcal{A}}\times\{t_{\infty}\}$. This will only be used in the next subsection.
@@
Let $\widetilde{{\mathcal{R}}}_{0}$ be the union of all irreducible components of $\widetilde{{\mathcal{R}}}$ which dominate some irreducible component of $\mathcal{R}$}. View $\widetilde{{\mathcal{R}}}_{0}$ as a $\mathbb{Q}$-Cartier divisor on $\widetilde{{\mathcal{X}}}$ such that for any irreducible component of $\widetilde{{\mathcal{R}}}_{0}$, its multiplicity in $\widetilde{{\mathcal{R}}}_{0}$ is exactly the same as its multiplicity in $\widetilde{{\mathcal{R}}}$. Denote $\widetilde{{\mathcal{R}}}_{1}=\widetilde{{\mathcal{R}}}-\widetilde{{\mathcal{R}}}_{0}$. Then $\widetilde{{\mathcal{R}}}_{1}$ is an effective $\mathbb{Q}$-Cartier divisor on $\widetilde{{\mathcal{X}}}$ whose image in $\mathcal{X}$} has codimension greater than 1.
@@
Let ${\mathcal{Z}}_{1}$ be the image of $\widetilde{{\mathcal{R}}}_{1}$ in ${\mathcal{A}}\times_{\mathbb{C}}T$. Use Raynaud–Gruson’s flattening theorem (cf. [RG71, Théorème 5.2.2]) to the composition ${\mathcal{Z}}_{1}\hookrightarrow{\mathcal{A}}\times_{\mathbb{C}}T\to T$, we get a proper birational map $g\colon T^{\prime}\to T$ such that the strict transform ${\mathcal{Z}}^{\prime}_{1}$ of ${\mathcal{Z}}_{1}$ under $g$ is flat over $T^{\prime}$. Note that ${\mathcal{Z}}^{\prime}_{1}$ is a closed subscheme of ${\mathcal{A}}\times_{\mathbb{C}}T^{\prime}$, and the inverse image of ${\mathcal{Z}}_{1}$ in ${\mathcal{A}}\times_{\mathbb{C}}T^{\prime}$ is the union of ${\mathcal{Z}}^{\prime}_{1}$ and some $T^{\prime}$-vertical closed subscheme ${\mathcal{V}}_{T^{\prime}}$ of ${\mathcal{A}}\times_{\mathbb{C}}T^{\prime}$ (i.e. the image of ${\mathcal{V}}_{T^{\prime}}$ in $T^{\prime}$ is not Zariski dense). After removing finitely many points, we can assume that $t_{n}$ has a unique pre-image $t^{\prime}_{n}$ in $T^{\prime}$ which does not lie in ${\rm pr}_{T^{\prime}}({\mathcal{V}}_{T^{\prime}})$ for all $n\geq 1$. Furthermore, we can assume $t^{\prime}_{n}$ tends to $t^{\prime}_{\infty}$ in the Euclidean topology. By our definition of $\widetilde{{\mathcal{R}}}_{1},{\mathcal{Z}}_{1},{\mathcal{Z}}^{\prime}_{1}$, we see that for any $t^{\prime}\in T^{\prime}$, $({\mathcal{Z}}^{\prime}_{1})_{t^{\prime}}$ has codimension greater than 1 in ${\mathcal{A}}\times\{t^{\prime}\}$. Let $Z_{1}$ be the image of $({\mathcal{Z}}^{\prime}_{1})_{t^{\prime}_{\infty}}\subseteq{\mathcal{A}}\times\{t^{\prime}_{\infty}\}$ in ${\mathcal{A}}\times\{t_{\infty}\}$. We will only use ${\mathcal{R}}_{0},{\mathcal{R}}_{1}$ (and other related notations such as $Z_{1}$) in the next subsection.
@@
Using Lemma 3.7 for the morphism ${\mathcal{D}}_{{\mathrm{red}}}\hookrightarrow{\mathcal{A}}\times_{\mathbb{C}}T\to T$ and the sequence $t_{n}\to t_{\infty}$ of points of $T$, we get a Zariski open dense subset ${\mathcal{D}}_{t_{\infty}}^{\circ}\subseteq{\mathcal{D}}_{t_{\infty}}$ satisfying the properties in Lemma 3.7. Furthermore, taking a subsequence of $\{t_{n}\}_{n\geq 1}$ if necessary, we may assume that $\{t_{n}\}_{n\geq 1}$ satisfies the conditions of remark 3.9. Denote ${\mathcal{D}}_{b,t_{\infty}}^{\circ}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}({\mathcal{D}}_{t_{\infty}}^{\circ})_{b}$.
@@
  2. (2)
    $S_{0}$ contains all the points $b\in B(\mathbb{C})$ such that $({\mathcal{D}}_{\rm red})_{b}$ or $(\widetilde{{\mathcal{R}}}_{\rm red})_{b}$ is non-reduced, which is finite by the generic reducedness (cf. [STA26, Lemma 054Z, Lemma 020I, Lemma 0578]).
  3. (3)
@@
Now we use Theorem 2.3 to construct a limit entire curve. Take $\ell_{n}\to\infty$ such that $\ell^{-1}_{n}s_{n}$ converges to a nonzero element $s_{\infty}$ in $V(A,K)$. Let $G=G(s_{\infty})$ be the smallest abelian subvariety of $A$ such that $s_{\infty}$ lies in $G(K)_{\mathbb{R}}$. Denote by $\mathcal{G}$} the Zariski closure of $G$ in $\mathcal{A}$}.
@@
As $b\notin S_{0}$, by property (6) of $S_{0}$, we can take an open disc ${\mathbb{D}}\subseteq B(\mathbb{C})$ with center $b$ such that $\mathbb{D}$} does not intersect with the image of ${\rm Supp}(\widetilde{{\mathcal{V}}}_{B})$ in $B$. After removing finitely many $n$, we may assume that $b_{n}\in{\mathbb{D}}$ for all $n\geq 1$. Apply Theorem 2.3 to $\{s_{n}\}_{n\geq 1}$ and the limit point $x\in({\mathcal{A}}_{b}\times\{t_{\infty}\})\backslash({\mathcal{D}}_{b,t_{\infty}}\cup{\mathcal{E}}^{\prime}_{b,t_{\infty}}\cup(Z_{1})_{b})$ of $\{s_{n}(B)\}_{n\geq 1}$. We obtain a linear entire curve $\varphi_{\infty}=\phi_{(x,\delta(s_{\infty},v_{b}))}\times\{t_{\infty}\}\colon\mathbb{C}\to{\mathcal{A}}_{b}\times\{t_{\infty}\}$, which is the unique limit of the re-parametrization $\{\varphi_{n}=\phi_{n}\times\{t_{n}\}\colon{\mathbb{D}}_{r_{n}}\to{\mathcal{A}}\times\{t_{n}\}\}_{n\geq 1}$ of $\{(s_{n},t_{n})\colon B\to{\mathcal{A}}\times\{t_{n}\}\}_{n\geq 1}$. Explicitly, recall that $\varphi_{n}$ is defined by
@@
For line bundle $L={\mathcal{O}}_{{\mathcal{A}}\times_{\mathbb{C}}T}({\mathcal{D}}_{\rm red})$ or $L={\mathcal{O}}_{{\mathcal{A}}\times_{\mathbb{C}}T}({\mathcal{E}}^{\prime})$ on ${\mathcal{A}}\times_{\mathbb{C}}T$ with the fixed metric $\left\|\cdot\right\|_{L}$ above, we choose the induced metric $(f\circ\pi)^{*}\left\|\cdot\right\|_{L}$ on $(f\circ\pi)^{*}L$. Moreover, for any $\mathbb{Z}$-linear combination of the line bundles above (including the two pullback line bundles), we can also fix a natural metric on this new line bundle induced by the metrics above.
For a line bundle $L$ which is a $\mathbb{Z}$-linear combination of the line bundles above (including the two pullback line bundles) with the fixed metric $\left\|\cdot\right\|_{L}$, a holomorphic map $\phi\colon{\mathbb{D}}_{R}\to\widetilde{{\mathcal{X}}}$ (or $\phi\colon{\mathbb{D}}_{R}\to{\mathcal{A}}\times_{\mathbb{C}}T$, depending on $L$ is a line bundle on $\widetilde{{\mathcal{X}}}$ or on ${\mathcal{A}}\times_{\mathbb{C}}T$), set
@@
Finally, the above definitions can be extended by $\mathbb{Q}$-linearity to any $\mathbb{Q}$-linear combination of the line bundles above (including the two pullback line bundles).
@@
Estimations on the limit entire curve $\varphi_{\infty}$
We first do some estimates on the limit entire curve $\varphi_{\infty}$. The main ingredient is Theorem 3.6(1). The philosophy is that, $N^{(1)}$ serves as the main part of the characteristic function $T\approx N^{(1)}+(N-N^{(1)})+m$, while $N-N^{(1)}$ and $m$ are negligible compared to $N^{(1)}$. This corresponds to inequalities (3.4.5) and (3.4.6) below, which will be used respectively when estimating the upper bound and the lower bound of $N-N^{(1)}$.
By assumption (2) for $x$, $({\mathcal{D}}_{b,t_{\infty}})_{\rm red}\cap(x+{\mathcal{G}}_{b}\times\{t_{\infty}\})$ is a reduced divisor on $(x+{\mathcal{G}}_{b}\times\{t_{\infty}\})$. Using Theorem 3.6(1) and the ampleness of ${\mathcal{D}}_{b,t_{\infty}}$ on ${\mathcal{A}}_{b}\times\{t_{\infty}\}$ (assumption (4) for $S_{0}$), we see that there exists a subset $\Sigma_{\varepsilon}\subseteq\mathbb{R}_{>0}$ of finite Lebesgue measure such that for any $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$,
@@
Therefore for any fixed $\varepsilon>0$ and any $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$, combine (3.4.2) and (3.4.4), we get
@@
Moreover, for fix $0<\varepsilon<\frac{1}{2}$ above, $\frac{\varepsilon}{1-\varepsilon}<2\varepsilon$ and $N^{(1)}(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})\geq\frac{c_{2}}{\varepsilon}$ when $r$ large enough. Thus by enlarging $\Sigma_{\varepsilon}$ we may assume that for all $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$,
@@
On the other hand, combine (3.4.2) and (3.4.3), we have for all $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$,